# Don’t Cut Corners: Exact Conditions for Modularity in Biologically Inspired Representations

### v1

@@ 4 Modularisation in Biologically Inspired Recurrent Neural Networks
4.1 Linear RNNs
Linear sinusoidal regression. Linear dynamical systems can only autonomously implement exponentially growing, decaying, or stable sinusoidal functions. We therefore study biologically inspired linear RNNs constrained to model stable sinusoidal signals at certain frequencies:
We study the optimal nonnegative, efficient, recurrent representations, ${\bm{z}}_{t}$, and show that whether the representations of the two frequencies within ${\bm{z}}_{t}$ modularises depends on their ratio. We prove and verify empirically that if one frequency is an integer multiple of the other the encodings mix, whereas if their ratio is irrational they should modularise. Further, we show empirically, and prove in limited settings, that rational non-harmonic ratios should modularise (Appendix C). The intuition for this result is much the same as the linear autoencoding setting: the natural notions of sources are the signals $(\cos(\omega_{1}t),\sin(\omega_{1}t),\cos(\omega_{2}t),\sin(\omega_{2}t))$. Using these sources, we must simply ask: does their support allow for a reduction in activity energy via mixing? In Figure 3a, we visualise the source support for the three prototypical relationships between $\omega_{1}$ and $\omega_{2}$: irrational, rational (but not harmonic), and harmonic. Results of neural network verifications are in Figure 3b (details Appendix G). In the irrational case, the source support is essentially rectangular, so the model modularises. In the harmonic case, large chunks of various corners are missing from the source support, so the model mixes. In the rational case, even though the source support is quite sparse, the corners are sufficiently present such that modularising is still optimal.
Modularisation of grid cells. We now show that these spectral ideas can explain modules of grid cells. Grid cells are neurons in the mammalian entorhinal cortex that fire in a hexagonal lattice of positions [22]. They come in groups, called modules; grid cells within the same module have receptive fields (firing patterns) that are translated copies of the same lattice, and different modules are defined by their different lattice [60]. Current theories suggest that the grid cell system can be modelled as a RNN with activations built from linear combinations of frequencies [14], just like the linear RNNs considered here. Importantly, these grid cell theories use the same biological constraints as in this work and show that modules form because the optimal code contains non-harmonically related frequencies that are encoded in different neurons (Figure 3e top). This modularisation can now be theoretically justified in our framework, since non-harmonic frequencies are range-independent and so should be modularised (Figure 3e bottom).
@@ 4.2 Nonlinear RNNs
Mixed sinusoidal regression for nonlinear RNNs. We train nonlinear ReLU RNNs with biologically inspired constraints to perform a frequency mixing task (details Appendix G). We provide a pulse input $P_{\omega}(t)=\mathbb{I}\left[\operatorname{mod}_{\omega}(t)=0\right]$ at two frequencies, and the network has to output the resulting “beats” and “carrier” signals:
@@
Modularisation in nonlinear teacher-student distillation. To test our predictions of when RNNs modularise, but in settings more realistic than pure frequencies, we generate training data trajectories from randomly initialised teacher RNNs with tanh activation function, and then train student RNNs (with a ReLU activation function) on these trajectories. The student’s representation is constrained to be nonnegative (via its ReLU) and has its activity and weights regularised (see Section G.2 for details). Using carefully chosen inputs at each timestep, we are able to precisely control the teacher RNN hidden state activity (i.e., the source distribution). This allows us to change correlations/statistical independence of the hidden states, while either maintaining or breaking range independence. We consider three settings (Figure 3e). First, when statistical and range independence are progressively broken (in orange). Second, where statistical, and to a lesser extent range, independence gets progressively broken (in green). Third, where only statistical independence gets progressively broken (in purple). We qualitatively observe, as our linear theory predicts, that students learns modular representations when range independence is maintained regardless of the statistical dependence of the teacher RNN hidden states (Figure 3f).
5 Modularisation of Prefrontal Working Memory
We now apply our results to neuroscience to explain a puzzling difference in monkey prefrontal neurons in two seemingly similar working memory tasks. In both tasks [71, 45], items are presented to an animal, and, after a delay, must be recalled according to the rules of the task. Similarly, in both tasks, as well as in neural networks trained to perform these tasks [8, 48, 70], the neural representation during the delay period consists of multiple subspaces, each of which encodes a memory of one of the items. Bizarrely, in one task these subspaces are orthogonal to one another [45], a form of modularisation in the representation of different memories sans a preferred basis, while in the other they are not [71].
In more detail, Xie et al. [71] train monkeys on a sequential working memory task in which the subject observes a sequence of $N$ dots positioned on a screen then, after a delay, must report that same sequence via saccading to the dot locations in order (Figure 4a). They find that the neural representation in the delay period decomposes into $N$ subspaces (one for each item) that are significantly non-orthogonal. (To calculate the alignment we reanalysed the data using a different measure of alignment from the original paper, inspired by Panichello and Buschman [45]. This is because we find other measures to be biased Appendix H). On the other hand, in a related task, Panichello and Buschman [45] (P&B) find orthogonal memory. They present monkeys with two coloured squares, one below the other, then, after a delay, present a cue that tells the monkey to recall the top or bottom colour via a saccade to the appropriate point on a colour wheel (Figure 4c). P&B find that, during the delay after cue presentation, the colours are encoded in two subspaces that are orthogonal to one another (Figure 4d).
Why does one experiment result in orthogonality but the other not? Our theory says that differences in range (in)dependence can lead to modular or mixed (orthogonal or not in this case) codes. Crucially, across trials, P&B sample the two colours independently, whereas Xie et al. [71] sample the dots without replacement. The latter sampling scheme compromises orthogonality since there are corners (indeed an entire diagonal) missing from the source support.
Figure 4: a) Xie et al. [71] task. b) We estimate the non-orthogonality of subspaces in data, and networks trained on sequences sampled with and without replacement or with correlations (Appendices H and I). Sampling without replacement, as in the experiment, gives the best fit. c) Panichello and Buschman [45] task, replicated. d) Subspace correlations for data and models. Sampling with replacement is necessary to fit the data.
We model this in biologically-constrained RNNs for each task. For the Xie et al. [71] task we present two inputs sequentially to a linear RNN with nonnegative hidden state and ask the networks to recall the inputs, in order, after a delay. Interestingly, our results depend on the particular choice of item encoding (one-hot or 2D), in ways that can be understood using our theory (Appendix I). To compare to the prefrontal data, we extract the item encoding from the prefrontal data during the delay period for single-element sequences. For P&B we assume the encoding is 2D.
Our simulated models recapitulate the major observations of monkey prefrontal representations. Sequences sampled without replacement in the Xie et al. [71] task lead to aligned encoding subspaces like data (Figure 4b), while colours sampled independently lead to orthogonal encodings in the P&B task as in data (Figure 4d). Conversely, as a prediction for future primate experiments, swapping the sampling scheme swaps the prediction. That said, there remain some puzzling discrepancies to explore. Panichello & Buschmann find that about 20% of colour-tuned neurons are tuned to the both colours, despite the orthogonal encoding [45], whereas our theory would predict all neurons would respond to one colour. Further, despite a qualitative match (i.e., orthogonal versus non-orthogonal), we do not obtain the exact angles between subspaces in Xie et al. using our RNNs (Figure 4b).
6 Orthogonalisation of Preparatory and Potent Subspaces in Motor Cortex
For our second neuroscience puzzle we turn to the motor cortex. Neural activity in primate motor cortex appears to drive movements, such as reaching. Further, in classic delayed-reach tasks, where the monkey is shown a cue then, after a delay, has to reach towards that cue, there is a delay-period activity that encodes which movement the monkey is preparing to make [33]. This has led researchers to break the motor cortical neural activity into two subspaces: a preparatory neural subspace, in which upcoming movements are prepared, and a potent subspace that actually drives behaviour. RNN models trained on delayed-reach tasks also show preparation related activity. However, while in neural data these two subspaces are orthogonal, in most RNN models the two are aligned [21]. Intriguingly, a further study trained RNNs to perform two cued-reaches rapidly, one after another, and found that then the preparatory and potent subspaces were orthogonal [75]. Why are these subspaces aligned in RNNs trained on single action tasks, but orthogonal in both the brain and networks trained on dual-reach tasks?
Figure 5: a) Dual-reach task. An RNN receives a cue that says which reach to perform (orange or blue), and a go signal. It must perform the required reach after the go cue. Network activity can be broken down into preparatory directions, which prepare the network to reach in a certain direction, and potent directions, which enact the movement. b) Following [21], we calculate an alignment index: the proportion of preparatory activity during single-reach trials explained by the potent subspace (details in Appendix I). For networks trained to perform only single reaches, the two subspaces are aligned; while for networks trained to perform dual-reach tasks, and neural data, the two are orthogonal. Finally, a network trained to perform two reaches, but in which the preparatory cue for the second movement arrives after the end of the first movement, again has aligned subspaces, matching our theory.
This can be understood in our framework as an independence effect. During the dual-reach, but not the single-reach, task, the preparatory and potent subspaces are both being concurrently used. This ensures that the activity in the two subspaces can be range independent from one another, whereas in the single-reach case either the preparatory or the potent signal is non-zero, but never both. Since independence drives orthogonalisation, this means that only in the dual-task case is there a push towards orthogonalising the representation. We train networks to perform simple analogues of the single- and dual-reach tasks (details Appendix I) and recover the same results as stated: only those trained on dual-reach tasks have close-to-orthogonal potent and preparatory subspaces Figure 5. Finally, this effect only occurs if preparation and movement overlap. If a network is trained on dual-reach tasks in which the second preparatory cue arrives after the end of the first movement, the subspaces again align Figure 5. Motor cortex likely orthogonalises these subspaces since it is regularly performing sequences of movements, preparing the next while performing another. This is natural, but does mean that our predictions are really about the behaviour of RNN models, since no motor cortex is likely to have range-dependent preparatory and potent activity over the course of the animal’s lifetime.
@@ 7 Modular or Mixed Codes for Space & Reward
Our last riddle comes from the entorhinal cortex. This brain area has been thought to contain precisely modular neurons, such as the grid cell code for self-position [22], object vector cells that fire at a particular displacement from objects [30], and heading direction cells [63]. Two recent papers examined the influence of rewarded locations on the grid cell code and find differing effects on the modularity of grid cells. Butler et al. [11] find that rewards rotate the grid cells, but preserve their pure spatial coding, while Boccara et al. [7] find that the grid cells warp towards the rewards, becoming mixed-selective to reward and space.
@@
However, with our improved understanding of modularisation, this makes sense. Despite Butler et al. [11] correlating reward and position, critically, they leave them range independent—all combinations of reward and position are possible. On the other hand, Boccara et al. [7] make certain combinations of reward and position impossible, making them not just correlated, but range correlated. As such, our theory matches this modularisation pattern. To test this we train a linear biological RNN (details Appendix I) to report both its self-position and displacement from a reward as it moves around an environment (Figure 6). We train RNNs on settings with different relationships between reward and position; for some RNNs the rewards are in fixed positions in every environment (range dependent and statistically dependent), for others reward and position are uniformly random in each room (range independent and statistically independent), and for a final group have both settings in different proportions (range independent but statistically dependent). As we vary the proportion of fixed rooms, we find the optimal representation stays modular, containing separate spatial and reward-vector cells as in Butler et al. [11], until all the sampled rooms have fixed reward positions, at which point neurons become mixed selective to reward and self-position, as in Boccara et al. [7]. As such, our theory now covers all known modularity results about grid cells.
8 The Mixed Origins of Mixed Selectivity
Why neurons might be mixed selective is a matter of debate in neuroscience [3, 65]. Theories of nonlinear mixed selectivity have argued that, analogously to a nonlinear kernel, such schemes permit linear readouts to decode nonlinear task functions of the sources. This is likely a key part of the mixed selectivity found in some brain areas, like the cerebellum [37], mushroom body [1], and perhaps certain prefrontal or hippocampal representations [6, 9]. However, our theory raises the possibility for other explanations of both nonlinear and linear mixed selectivity that do not require tasks including nonlinear functions of the sources.
Figure 7: a, b) We train linear RNNs to report displacement to three objects as an agent moves within many rooms. In each room, object instances are either completely random (a, top) or clustered in a random line (a, bottom); object positions are therefore range independent but correlated. c) The neurons are modular: each neuron’s activity is conditionally informative of only a single object given the others. d) However, if an experimenter were only aware of two of the three objects, the neurons that purely encode the disregarded object would appear mixed selective due to the statistical dependencies between objects. e) We extract the tuning of neurons to different memory encodings from both RNNs (left) and neural data (right) [71]. In both the models and data, many neurons are linearly mixed selective. f) In nonlinear, but linearly decoded, representations from the What-Where Regression task, many neurons are nonlinearly mixed selective. g) Warden and Miller [67] train monkeys to see, then report a sequence of two images. Mikkelsen et al. [41] reanalyse the data and find h) neurons tuned to both time and stimulus identity. i) Our linear RNN model recapitulates this.
@@ 8 The Mixed Origins of Mixed Selectivity
Missing sources. First, a purportedly mixed selective neuron could actually be a modular encoding, but of a source variable unknown to the experimenter. For example, despite the many modular neurons in entorhinal cortex, there are also many neurons that seem mixed selective to combinations of spatial variables, such as speed, heading direction, or position [23]. An alternate explanation is that these neurons may (in part) be purely selective for another unanalysed variable that is itself correlated with the measured spatial variables or behaviour. We highlight a simple example of this effect in Figure 7a, b. Imagine a mouse is keeping track of three objects as it moves in an environment (Figure 7a); we model this as a linear biological RNN that must report the displacement of all objects from itself (Figure 7b). Object positions are range independent but correlated, so the RNN representation modularises such that each neuron only encodes one object (Figure 7c). However, if an experimenter were only aware of two of the three objects, they would instead analyse the neural tuning with respect to those two objects. Due to the statistical dependencies between object positions, they would find mixed selective neurons that are in reality purely coding for the missing object (Figure 7d).
Range-dependent sources. Second, our theory gives a precise set of inequalities for (biologically constrained) modularisation of scalar sources. Breaking the condition means the optimal representation is linearly mixed selective (Figures 1, 3, 4 and 7). These ideas also qualitatively predict modularisation of nonlinear feedforward and recurrent networks across a range of settings (Figures 2 and 3). Figure 7e, f re-illustrates these results, which suggest that nonlinear mixed selectivity might exist simply to save energy, rather than the typical rationale of permitting flexible linear readouts of arbitrary categorisations [3, 65].
Sequential processing. Finally, multiple works have highlighted nonlinear mixed selectivity to stimulus feature and time within task [46, 12, 13]. We show that simple linear sequential computations can produce this seemingly nonlinear temporal mixed selectivity. For example, Warden and Miller [67] train monkeys to view sequences of two images separated fixed delays (Figure 7g). To be rewarded, monkeys must report the stimuli, e.g., by saccading to the images in the correct sequence. Recently, Mikkelsen et al. [41] reanalysed this data and found that many neurons are nonlinearly mixed selective to stimulus and time (Figure 7h). However, we train a biological linear RNN to recall two sequentially presented one-hot cues separated by delays, and find similar neurons (Figure 7i and Appendix I). This form of nonlinear mixed selectivity arises not from downstream readout pressures, but simply as the optimal energy-efficient code during fixed-length delays. Indeed, interpreting these as nonlinear relies on one viewing time as a scalar linearly increasing variable; if instead time is represented one-hot, both monkey and model neurons could be interpreted as linearly mixed selective.
@@ 9 Discussion
In this work we have given precise constraints that cause linear feedforward and recurrent networks under biological constraints to modularise. These constraints are highly dependent on the co-range properties of variables, and have helped us understand patterns of modularisation in both nonlinear networks and the brain.
Theoretical limitations. Our theory predicts binary yes/no modularity in linear sources (each uni-dimensional) with L2 energy constraints, but does not yet predict the modularisation of multidimensional sources, nonlinear settings, other norms (using an L1 sparsity penalty on neural activity has a long history [44]), nor does it prescribe the degree of modularity (rather than perfectly modular or not). These could all be useful extensions.
Mixed selectivity. In neuroscience our findings add nuance to the ongoing debate over mixed vs. modular neural coding. Our theory points to a variety of settings in which mixed-selectivity arises under energy constraints, without the requirement of a flexible linear readout for nonlinear classifications that is often argued to the the source of mixed-selectivity. Nevertheless we think that flexible linear readouts are sometimes the source of mixed-selectivity, to allow kernel regression-like classification. We expect this is true for neurons in the cerebellum or mushroom body, as well as the some cortical neurons that have non-linearly mixed selective coding of independent variables—something our theory would never predict [6]. Yet our theory suggests several alternative explanations for mixed-selectivity that seem plausible for many recorded neural populations, for example those in entorhinal (Figure 6) and prefrontal cortex (Figure 4) including the dataset for which these nonlinear classification theories were originally developed [67].
Neural limitations. Our theory prescribes when biological neural populations should be modular or not, yet some of the PFC (Section 5) and motor cortex (Section 6) data we explain is when the population is orthogonal or not. Orthogonal is not necessarily modular; indeed some real individual neurons are seemingly not fully modular, contrary to our theory. This may be a shortcoming of our simple theory. However, it also possible that it is best to understand these mixed-selective neurons, for example in Panichello and Buschman [45]’s task, as a different computational population. This feature—pure and mixed selective cells playing different computational roles—is observed in entorhinal cortex: pure grid cells code only for spatial position while conjunctive grid/heading-direction cells [54] are thought to be critical in integrating velocity to update the position code [66] (much like the analogous heading-direction update neurons in the protocerebral bridge of the fly central complex [64]). We anticipate that incorporating computational elements into our theory will further account for mixed selective neurons in frontal cortex and beyond.
@@
Disentangling what? In machine learning, our work bolsters the importance of range properties over statistical independence in the disentanglement literature. On this we agree with Roth et al. [52], who argue that range independence is the more meaningful form of independence. Further, our conditions have qualitative similarities to sufficient spread conditions from the matrix factorisation literature [61, 62], perhaps suggesting a deeper correspondence. It is intriguing that the constraints inspired by those in biology, which have shown promise for disentangling [69], and have links to state-of-the-art disentangling methods [27, 28], are sensitive to just this form of independence.
Acknowledgements
The authors thank Basile Confavreux and Pierre Glaser for conversations about proof techniques.
We thank the following funding sources: Gatsby Charitable Foundation to W.D., J.H.L. & P.E.L.; Sequoia Capital Stanford Graduate Fellowship and NSERC Postgraduate Scholarship – Doctoral to K.H.; Wellcome Trust (110114/Z/15/Z) to P.E.L.; Wellcome Trust (219627/Z/19/Z) to J.H.L.; Sir Henry Wellcome Postdoctoral Fellowship (222817/Z/21/Z) to JCRW.; Wellcome Principal Research Fellowship (219525/Z/19/Z), Wellcome Collaborator award (214314/Z/18/Z), and Jean-François and Marie-Laure de Clermont-Tonnerre Foundation award (JSMF220020372) to TEJB.; the Wellcome Centre for Integrative Neuroimaging is supported by core funding from the Wellcome Trust (203139/Z/16/Z).
Reproducibility
Code used in producing this work is available at https://github.com/kylehkhsu/dont-cut-corners/.

### v2

@@ 4 Modularisation in Biologically Inspired Recurrent Neural Networks
4.1 Linear RNNs
Linear sinusoidal regression. Linear dynamical systems can only autonomously implement exponentially growing, decaying, or stable sinusoidal functions. We therefore study linear RNNs with biological constraints trained to model stable sinusoidal signals at certain frequencies:
We study the optimal nonnegative, efficient, recurrent representations, ${\bm{z}}_{t}$, and show that whether the representations of the two frequencies within ${\bm{z}}_{t}$ modularises depends on their ratio. We prove and verify empirically that if one frequency is an integer multiple of the other the encodings mix, whereas if their ratio is irrational they should modularise. Further, we show empirically, and prove in limited settings, that rational non-harmonic ratios should modularise (Appendix E). The intuition for this result is much the same as the linear autoencoding setting: the natural notions of sources are the signals $(\cos(\omega_{1}t),\sin(\omega_{1}t),\cos(\omega_{2}t),\sin(\omega_{2}t))$. Using these sources, we must simply ask: does their support allow for a reduction in activity energy via mixing? In Figure 3a, we visualise the source support for the three prototypical relationships between $\omega_{1}$ and $\omega_{2}$: irrational, rational (but not harmonic), and harmonic. Results of neural network verifications are in Figure 3b (details Appendix I). In the irrational case, the source support is essentially rectangular, so the model modularises. In the harmonic case, large chunks of various corners are missing from the source support, so the model mixes. In the rational case, even though the source support is quite sparse, the corners are sufficiently present such that modularising is still optimal.
Modularisation of grid cells. We now show that these spectral ideas can explain modules of grid cells (literature review: Appendix A. Grid cells are neurons in the mammalian entorhinal cortex that fire in a hexagonal lattice of positions [17]. They come in groups, called modules; grid cells within the same module have receptive fields (firing patterns) that are translated copies of the same lattice, and different modules are defined by their different lattice [48]. Current theories suggest that the grid cell system can be modelled as a RNN with activations built from linear combinations of frequencies [11], just like the linear RNNs considered here. Importantly, these grid cell theories use the same biological constraints as in this work and show that modules form because the optimal code contains non-harmonically related frequencies that are encoded in different neurons (Figure 3e top). This modularisation can now be theoretically justified in our framework, since non-harmonic frequencies are range-independent and so should be modularised (Figure 3e bottom).
@@ 4.2 Nonlinear RNNs
Mixed sinusoidal regression for nonlinear RNNs. We train nonlinear ReLU RNNs with biologically inspired constraints to perform a frequency mixing task (details Appendix I). We provide a pulse input $P_{\omega}(t)=\mathbb{I}\left[\operatorname{mod}_{\omega}(t)=0\right]$ at two frequencies, and the network has to output the resulting “beats” and “carrier” signals:
@@
Modularisation in nonlinear teacher-student distillation. To test our predictions of when RNNs modularise, but in settings more realistic than pure frequencies, we generate training data trajectories from randomly initialised teacher RNNs with tanh activation function, and then train student RNNs (with a ReLU activation function) on these trajectories. The student’s representation is constrained to be nonnegative (via its ReLU) and has its activity and weights regularised (see Section I.2 for details). Using carefully chosen inputs at each timestep, we are able to precisely control the teacher RNN hidden state activity (i.e., the source distribution). This allows us to change correlations/statistical independence of the hidden states, while either maintaining or breaking range independence. We consider three settings (Figure 3e). First, when statistical and range independence are progressively broken (in orange). Second, where statistical, and to a lesser extent range, independence gets progressively broken (in green). Third, where only statistical independence gets progressively broken (in purple). We qualitatively observe, as our linear theory predicts, that students learns modular representations when range independence is maintained regardless of the statistical dependence of the teacher RNN hidden states (Figure 3f).
@@ 5 Modular or Mixed Codes for Space & Reward
We now apply our results to neuroscience to understand a puzzling difference in two seemingly similar recordings from entorhinal cortex. This brain area has been thought to contain precisely modular neurons, such as the grid cell code for self-position [17], object vector cells that fire at a particular displacement from objects [25], and heading direction cells [51]. Two recent papers examined the influence of rewarded locations on the grid cell code and find differing effects on the modularity of grid cells. Butler et al. [10] find that rewards rotate the grid cells, but preserve their pure spatial coding, while Boccara et al. [6] find that the grid cells warp towards the rewards, becoming mixed-selective to reward and space.
@@
However, with our improved understanding of modularisation, this makes sense. Despite Butler et al. [10] correlating reward and position, critically, they leave them range independent—all combinations of reward and position are possible. On the other hand, Boccara et al. [6] make certain combinations of reward and position impossible, making them not just correlated, but range correlated. As such, our theory matches this modularisation pattern. To test this we train a linear RNN with biological constraints (details Appendix J) to report both its self-position and displacement from a reward as it moves around an environment (Figure 4). We train RNNs on settings with different relationships between reward and position; for some RNNs the rewards are in fixed positions in every environment (range dependent and statistically dependent), for others reward and position are uniformly random in each room (range independent and statistically independent), and for a final group have both settings in different proportions (range independent but statistically dependent). As we vary the proportion of fixed rooms, we find the optimal representation stays modular, containing separate spatial and reward-vector cells as in Butler et al. [10], until all the sampled rooms have fixed reward positions, at which point neurons become mixed selective to reward and self-position, as in Boccara et al. [6]. As such, our theory now covers all known modularity results about grid cells.
@@
Missed sources and mixed selectivity. In contrast to the beautifully modular neurons in entorhinal cortex, recent work has highlighted neurons with mixed tuning to multiple navigational variables, such as position, heading direction, or speed [18]. We use our theory to highlight a potential artifact of this type of analysis. A purportedly mixed selective neuron may (in part) be purely selective for another unanalysed variable that is itself correlated with the measured spatial variables or behaviour. We highlight a simple example of this effect in Figure 5a, b. Imagine a mouse is keeping track of three objects as it moves in an environment (Figure 5a); we model this as a linear RNN with biological constraints that must report the displacement of all objects from itself (Figure 5b). We make the object positions range independent but correlated. As predicted by our theory, the RNN modularises these range-independent variables such that each neuron only encodes one object (Figure 5c). However, if an experimenter were only aware of two of the three objects, they would instead analyse the neural tuning with respect to those two objects. Due to the statistical dependencies between object positions, they would find mixed selective neurons that are in reality purely coding for the missing object (Figure 5d). We do not believe that all neurons in entorhinal cortex are modular; for example, there is clear evidence that neurons co-tuned to position and velocity are meaningfully mixed-selective, implementing the path-integration updates in the grid cell system [53]. Indeed, many of the neurons analysed in [18] likely are mixed selective. However, it is also seems likely that in more exploratory analyses such as these, mixed selectivity might often arise from missing encoded variables in the analysis.
Figure 5: a, b) We train linear RNNs to report displacement to three objects as an agent moves within many rooms. In each room, object instances are either completely random (a, top) or clustered in a random line (a, bottom); object positions are therefore range independent but correlated. c) The neurons are modular: each neuron’s activity is conditionally informative of only a single object given the others. d) However, if an experimenter were only aware of two of the three objects, the neurons that purely encode the disregarded object would appear mixed selective due to the statistical dependencies between objects.
@@ 6 Discussion
We present precise constraints on source distributions that cause linear feedforward and recurrent networks with biological constraints to modularise. These constraints depend on the co-range properties of variables, matching patterns of modularisation in both nonlinear networks and the brain.
Limitations & future work. That said, there remain many aspects of our theory that need improvement. Deriving necessary (rather than sufficient) modularisation conditions for other norms and of multidimensional sources, extending to nonlinear settings, or to understand a less binarised notion of modularity (rather than perfectly modular or not) are attractive directions. Additionally, our model is a purposefully simple model, there are many biologically pertinent effects whose impact on modularity would be interesting to study, such as connection sparsity, anatomical constraints, noise, or requiring additional computational roles from the network. Finally, the theory needs testing beyond entorhinal cortex.
Mixed selectivity. In neuroscience our findings add nuance to the ongoing debate over mixed vs. modular neural coding. Theories of nonlinear mixed selectivity have argued that, analogously to a nonlinear kernel, such schemes permit linear readouts to decode nonlinear functions of the sources [39]. This is likely a key part of the mixed selectivity found in some brain areas, like the cerebellum [31], mushroom body [1], and perhaps certain prefrontal or hippocampal representations [5, 7]. However, our theory raises the possibility for other explanations of both nonlinear and linear mixed selectivity that do not require tasks with nonlinear functions of the sources. First, purely from energy efficiency, our work analytically shows when range-dependent variables should be encoded in linearly mixed-selective representations, and empirically shows the same ideas apply to nonlinear mixed-selective representations Figure 2. Further, since in our theory even statistically dependent variables might be represented modularly, we show that this causes problems for analysis techniques that do not account for all the variables encoded in a population, leading to spurious mixed selectivity Section 5. Future work could usefully developed principled approaches to distinguishing these options.
@@
Disentangling what? In machine learning, our work bolsters the importance of range over statistical independence. We agree with Roth et al. [41], who argue that range independence is the more meaningful form of independence. Further, our conditions have qualitatively similarities to sufficient spread conditions from the matrix factorisation literature [49, 50], suggesting a deeper correspondence. It is intriguing that the constraints inspired by biology, which hold promise for disentangling [56], and have links to state-of-the-art disentangling methods [22, 23], are sensitive to precisely this form of independence.
Reproducibility: For code to reproduce the figures shown here please consult the following github repository: https://github.com/kylehkhsu/modular.